Suppose that the New England Colonials baseball team is equally likely to win a game as not to win it. If 5 Colonials games are chosen at random, what is the probability that exactly 3 of those games are won by the Colonials?
Round your response to at least three decimal places. (If necessary, consult a list of formulas.)


Answer by Boreal(15235)

This is 5C3*.5^3*.5^2=10*(0.03125)=0.3125
binomial nCr*p^r(1-p)^(n-r)